# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ Abstract
Epithelial tissue elongation by convergent extension is a key motif of animal morphogenesis. On a coarse scale, cell motion resembles laminar fluid flow; yet in contrast to a fluid, epithelial cells adhere to each other and maintain the tissue layer under actively generated internal tension. To resolve this apparent paradox, we formulate a model in which tissue flow in the tension-dominated regime occurs through adiabatic remodeling of force balance in the network of adherens junctions. We propose that the slow dynamics within the manifold of force-balanced configurations is driven by positive feedback on myosin-generated cytoskeletal tension. Shifting force balance within a tension network causes active cell rearrangements (T1 transitions) resulting in net tissue deformation oriented by initial tension anisotropy. Strikingly, we find that the total extent of tissue deformation depends on the initial cellular packing order. T1s degrade this order so that tissue flow is self-limiting. We explain these findings by showing that coordination of T1s depends on coherence in local tension configurations, quantified by a geometric order parameter in tension space. Our model reproduces the salient tissue- and cell-scale features of germ band elongation during Drosophila gastrulation, in particular the slowdown of tissue flow after approximately twofold elongation concomitant with a loss of order in tension configurations. This suggests local cell geometry contains morphogenetic information and yields experimentally testable predictions. Furthermore, defining biologically controlled active tension dynamics on the manifold of force-balanced states may provide a general approach to the description of morphogenetic flow.
NHC and FB contributed equally.
  Significance. Animal development requires large numbers of cells to choreograph their force generation in order to sculpt tissues and organs. Leveraging the fact that cellular forces equilibrate rapidly compared to the speed of development, we formulate a geometrical model for the network of balanced active tensions in an epithelial sheet. Within this framework, we can investigate how cells remodel the tension network to change tissue shape. A simple “winner-takes-all” mechanical feedback loop can self-organize complex cell movement, matching experimental data on the cell and tissue scale. We find that the ability to self-organize depends on initial order in the cellular packing. Our model explains how genetic patterning, embryo geometry, and cellular packing geometry combine to determine tissue shape change.
@@
Shape changes of epithelia during animal development involve major cell rearrangements, often manifested as a “convergent extension” of cell sheets. On the coarse scale, convergent extension looks much like the laminar shear flow of an incompressible fluid in the vicinity of a hyperbolic fixed point (see Fig. 1A). Indeed, previous work has combined hydrodynamic equations for the mesoscale cell velocity field with active stress fields to model morphogenetic tissue flow [1, 2, 3, 4]. Yet in contrast to a fluid, epithelia are under internally generated tension – revealed by laser ablation [5] – and, like solids, maintain their shape against external forces. Tissue flow is achieved through local cell intercalation (T1 neighbor exchange processes, see Fig. 1B) driven by the concerted mechanical activity of individual cells. Cells generate forces via actomyosin contractility in the cortical cytoskeleton at the adherens junctions between cells (Fig. 1C). Moreover, the adherens junctions can remodel through the turnover of its constituent molecules. Taken together, this implies that interfaces in the cell array can change their length and tension independently. This behavior is fundamentally different from (Hookean) springs, where tension and length are related by a constitutive relationship. Instead, one can imagine cellular interfaces as “microscopic muscles” which are controlled by the recruitment and release of myosin motors.
@@
Vertex models generally describe epithelial tissue as a polygonal tiling of cells where the vertex positions are the dynamical variables [6, 7]. The forces that drive the vertex motion are commonly derived from passive area and perimeter elasticity supplemented with additional active tensions [8, 9, 10]. However, the muscle metaphor for cellular interfaces suggests that active tension is central to the mechanical network underlying an epithelial tissue (Fig. 1D). This network rapidly equilibrates to a force balanced state [11, 5, 12], stabilized by mechanical feedback loops [13, 14]. In such an active tension network, passive elasticity plays a subdominant role. The need for stabilizing feedback loops arises because active tensions are untethered from interface lengths. Indeed, on an abstract level, these feedback loops are not unlike the regulatory mechanisms that control and stabilize skeletal musculature [15].
@@
Here, we propose that tissue flow can be understood in the terms of an adiabatic remodeling of internal active force balance. Force balance in the cortical tension network defines a manifold of cellular tiling geometries on which tissue deformation unfolds. We propose that dynamics in the force-balance manifold is driven by positive feedback on the cortical tensions. This view is supported by analysis of high quality live imaging data [16] from Drosophila gastrulation presented in the companion paper [17]. Specifically, a geometric active–passive decomposition as well as mutant analysis were used to show that tissue flow is driven by internally generated tension dynamics, rather than external forces. Tension inference has provided evidence for the role of positive tension feedback during active T1 events. Numerical simulations of cell quartets show that such a feedback mechanism is sufficient to drive the T1 process. However, the question of coordination of T1s across the tissue – required to drive coherent tissue flow – has remained unanswered. To address this question, we develop a model of tissue mechanics in the tension-dominated regime and demonstrate via numerical simulations how positive feedback drives convergent extension. We show that order of the cell packing is necessary for coordinating T1 processes, and hence efficient convergent extension. T1s destroy this order such that the extent of tissue flow is self limiting. Thereby, our model reproduces the experimentally observed elongation of the germband where the arrest of flow is concomitant with a transition from an ordered to a disordered cell packing [17].
@@ Methods
Our model is based on the assumption that on the timescales relevant to morphogenetic dynamics, the dominant active forces in the epithelium (generated by contractile myosin-2 motors in the adherens junctions, see Fig. 1C) are approximately balanced. In particular, we assume that adhesion forces between the epithelial layer and its substrate (the fluid yolk and perivitelline fluid [18] in the case of the Drosophila embryo) caused by the relatively slow morphogenetic motion, are negligible. Hence, all forces must be balanced within the transcellular network of cellular junctions. We model the tissue in the general framework of vertex models (see e.g. [7]) as a polygonal tiling of the plane with tri-cellular vertices $\mathbf{r}_{ijk}$, where each polygon represents a cell $i$ (see Fig 1D). We focus on a setting where active cortical tensions dominate over passive elasticity. We write the elastic energy differential of this network as
@@
where $\varepsilon$ is a small parameter that separates the scale of active and passive mechanical contributions. $E_{\mathcal{C}}(S_{i})$ accounts for the passive elasticity of the cells and will be specified below; $\ell_{ij}=||\mathbf{r}_{ij}||$ is the length of the interface between adjacent cells $i$ and $j$ and $T_{ij}$ is the cortical tension at that interface. Importantly, in contrast with the standard vertex model, where edge tension is defined by a constitutive relation corresponding to a passive Hookean perimeter spring, we take cortical tensions to be controlled independently of the interface lengths. The tension dynamics is described in the next section. The second term in Eq. (1) accounts for the effective in-plane pressure $p$ of the cells that, by maintaining the total surface area (sum over individual cell areas $A_{i}$), ensures that the tissue as a whole does not collapse. In the Drosophila embryo, this is a consequence of the fact that the closed epithelial sheet encloses the yolk, which is under pressure, balancing the cortical tensions. Other epithelia are kept under tension via traction forces at the boundary [19]. We assume that pressure differences between cells are small and therefore absorb them into $E_{\mathcal{C}}(S_{i})$. In experimental data, the assumption of small pressure differentials can be verified by inspecting the curvature of cell-cell interfaces: Due to the Laplace law, significant forces due to pressure differences would manifest through curved cell-cell interfaces which are not observed in the early Drosophila embryo prior to the onset of cell divisions [17, 16, 20, 12].
We further assume a separation of scales between the timescale on which the elastic energy relaxes and the timescale on which the tissue deforms macroscopically. In other words, we assume relaxational dynamics $\gamma\partial_{t}\mathbf{r}_{ijk}=-\frac{\partial E}{\partial\mathbf{r}_{ijk}}$, with a relaxation rate – set by the coefficient of friction $\gamma$ – much faster than all other timescales in the system. (Rather than frictional dissipation, one could equally use viscous dissipation here). Quasi-static force balance then implies
@@
Solving this equation to zeroth order in $\varepsilon$ yields a force-balance constraints at each vertex. These constraints imply that the tension vectors $\mathbf{T}_{ij}=T_{ij}\mathbf{e}_{ij}$ at each vertex sum to zero and hence form a triangle as illustrated in 1D. Since neighboring vertices share the interface that connects them, the corresponding tension triangles share an edge. Therefore, all tension triangles have to fit together to tile the plane: they form a triangulation that is dual to the cell tiling [13, 21, 22]. This tension triangulation is a geometric manifestation of global force balance in the tissue, where angles at real-space vertices are complementary to the corresponding angles in the tension triangle angles. One pair of such angles is highlighted in Fig. 1D.
Importantly, fixing the angles at vertices does not fully determine the cell tessellation, i.e. the $\mathbf{r}_{ijk}$ remain underdetermined, as this leaves the freedom to change the interface lengths $\ell_{ij}$ while preserving all angles. The resulting isogonal soft modes [^1] A simple counting of degrees of freedom shows that there is one isogonal degree of freedom per cell [13]. Therefore, the isogonal modes can be parametrized by an “isogonal function” that takes a scalar value in each cell. The isogonal displacement of a vertex is defined in terms of the values of this isogonal function in the three adjacent cells (see Eq. (S7)). thus account for interface length changes under constant tension [13], which is possible thanks to the turnover of cytoskeletal elements.
The isogonal modes can dilate and shear cells as illustrated in Fig. 1F. They are soft modes of the leading order force balance constraints and thus force us to take into account subleading contributions to the elastic energy to arrive at a complete model of tissue mechanics. Cells resist shape distortions due to rigid cell-internal structures such as microtubules, the nucleus [23, 24], and intermediate filaments [25]. To account for this passive cell elasticity, we propose an energy
@@
where $\mathcal{N}_{i}$ is the set neighbors of cell $i$. This shape energy effectively models the cell interior as a homogeneous elastic material [26]. (In SI Sec. 4, we compare this elastic energy with the often-used “area–perimeter” elastic energy which we find produces qualitatively different behavior that is incompatible with experimental observations.) The shape tensor is defined so that it is invariant under subdivision of interfaces. The reference tensor $S_{0}$ controls the target cell shape and is given by $S_{0}=3\ell_{0}\,\mathbb{I}$ for an isotropic hexagonal cell with side length $\ell_{0}$. External forces will deform the cell tessellation by acting on the isogonal modes. We can therefore relate cell elasticity and tissue elasticity by analyzing the energy spectrum of isogonal modes for a given tension triangulation (see SI Sec. 3). The two isogonal modes with the lowest energy correspond to uniform shear (see Fig. S13A). From these eigenmodes, we find a linear relationship between the cell and tissue shear moduli (see Fig. S13B).
@@
For sufficiently small values of the scale-separation parameter $\varepsilon$, minimization of the elastic energy Eq. (1) can be performed in two separate steps: First, force balance of the dominant cortical tensions $T_{ij}$ fixes the angles at vertices thus setting geometric constraints, but leaving isogonal modes undetermined. Second, the sub-leading term $\sum_{i}E_{\mathcal{C}}(S_{i})$ is minimized to fix these soft modes. Importantly, in this limiting case, the value of $\varepsilon$ is immaterial as long as it is small enough (see SI Sec. 2). Figure 1G and Movie 1 illustrate the minimization of the cell shape energy under the angle constraints imposed by junctional force balance for the minimal setting of a symmetric, regular cell array. The geometry is characterized by two interface lengths, $\ell_{0}$, $\ell_{1}$, and single angle $\phi$. Incompressibility fixes $\ell_{1}$ as a function of $\ell_{0}$ and $\phi$ such that only a single isogonal degree of freedom $\ell_{0}$ remains. We can plot the cell shape energy $E_{\mathcal{C}}$ in a two-dimensional energy landscape with the horizontal axis representing the tension-imposed angle constraint and the vertical axis representing the isogonal degrees of freedom. Force balance of the tensions $T_{0}$, $T_{1}$ constrains the angle $\phi=2\arccos(T_{0}/2T_{1}))$ and relaxation of $E_{\mathcal{C}}$ takes place in the isogonal subspace (illustrated by the vertical white line in Fig. 1G), thus fully determining the cell geometry (purple dot). Changing the tensions $T_{0},T_{1}$ shifts the angle constraint and therefore forces the cell shape energy to relax to a new cell geometry. This changing of constraints mechanism is how dynamic remodeling of the tension configuration drives tissue flow. When the tensions reach the critical ratio $T_{0}/T_{1}=\sqrt{2}$ such that $\phi=\pi/2$, the length of the vertical interface, $\ell_{0}$, vanishes causing a T1 transition as discussed in the companion paper [17]. To complete the model, we need to specify the dynamics governing the active tensions.
Figure 1: A model for tissue mechanics dominated by cortical tension. A Convergent extension of epithelial tissue by cell intercalations (T1 processes, B). C A single cell from a columnar epithelium with the actomyosin cortex at the adherens junction belt generating tension $T$ along cell-cell interfaces. Gray sphere represents the nucleus. D In force balance, the tensile forces $\mathbf{T}_{ij}$ at each vertex must sum to zero, thus linking the tensions to the angles at which the interfaces meet. Force balance implies that the tensile force vectors at each vertex form a triangle. Triangles at neighboring vertices share an edge and thus form a triangulation that tiles the plane (i.e. they have to fit together without gaps or overlaps). This defines a “tension space” that is dual to physical space. The angles in the tension triangulation are complimentary to the angles in the cell array, thus linking tension space to physical space. E Outline of our tissue model with mechanics dominated by actively regulated cortical tensions. Tension dynamics in adiabatic force balance are formulated as a geometric model of the tension triangulation which determines the angles at vertices. The remaining soft modes defineisogonal (angle-preserving) degrees of freedom that are governed by cell-shape elastic energy, which accounts for (passive) non-cortical elasticity and penalizes dilation and shearing. F Illustration of isogonal deformations of a reference geometry (purple) that is dual to the tension triangulation (red). G Implementation of the model for a symmetric, regular cell array, characterized by one angle $\phi$, determined by the tensions $T_{0},T_{1}$, and two lengths, $\ell_{0}$ and $\ell_{1}$, parametrizing the soft isogonal modes. The contour plot shows the cell shape energy $E_{\mathcal{C}}$ in the incompressible limit where $\ell_{1}$ is determined uniquely by $\ell_{0}$ and $\phi$. Relaxation of the sub-dominant cell shape energy $E_{\mathcal{C}}$ is constrained to the isogonal subspace (white line) determined by tension force balance. For the critical tension ratio $T_{0}/T_{1}=\sqrt{2}$, constraining the angle to $\phi_{c}=\pi/2$ (black dashed line), the interface length minimizing $E_{\mathcal{C}}$ vanishes (red half-disk). See Movie 1 for an animated version.
@@
Based on previous experiments [27] and models [28, 10], we propose a positive feedback mechanism where tension leads to further recruitment of myosin and thus further increase in tension. This self-amplifying accumulation of myosin on individual interfaces is limited by the competition for a limited pool of myosin within each cell. To mimic this local competition for myosin in a computationally simple way, we constrain tension dynamics to conserve the perimeter of each tension triangle, i.e. the sum of tensions at each vertex $(ijk)$. For an individual triangle with tensions $\tilde{T}_{1},\tilde{T}_{2},\tilde{T}_{3}$, we consider the dynamics
@@
where $n$ is an exponent that determines the nonlinearity of the feedback. Note that each cell-cell interface is composed of two cortices on the two sides of the interface, and only one of the two is part of each local pool (see SI Sec. 4 for details). This feedback mechanism has a “winner-takes-all” character, where the longest edge in the tension triangle always outgrows the other two. In our model framework we can consider a variety of possible local tension dynamics. Below we will also investigate a form of positive tension feedback that saturates, and identify the qualitative features of local tension dynamics key to the tissue dynamics.
@@
Force balance requires that all tension triangles fit together to form a flat triangulation [13]. The triangulation is parameterized by a set of 2D tension vertex positions $\mathbf{t}_{i}$, so that the tension on edge $(ij)$ is given by $T_{ij}=||\mathbf{t}_{i}-\mathbf{t}_{j}||$. In each iteration of the simulation, the tension vertices $\mathbf{t}_{i}$ are determined by fitting the balanced tensions $T_{ij}$ to the intrinsic tensions $\tilde{T}_{ij}$ using a least squares method. In addition, the intrinsic tensions $\tilde{T}_{ij}$ relax to the balanced tensions $T_{ij}$ with a rate $\tau_{\mathrm{balance}}^{-1}$ (see SI Sec. 4 for details). This “balancing” of the tension triangulation effectively accounts for small pressure differentials and additional feedback mechanisms (such as the strain rate feedback [13, 14]) which maintain the tension network in a state of force balance. In particular, this ensures that cortical tensions do not lead to build-up of pressure differentials. The rate at which these mechanisms act is controlled by the timescale $\tau_{\mathrm{balance}}$ which we assume to be much smaller than the timescale $\tau_{\mathrm{T}}$ on which tension evolve due to positive feedback. All quantifications presented here refer to the flat tensions $T_{ij}$.
@@
The above dynamics is autonomous in tension space until an edge in the cell tessellation reaches length zero. At this point, a cell neighbor exchange (T1 transition) occurs, corresponding to an edge flip in the tension triangulation. After this topological modification, the tension dynamics continues autonomously again until the next T1 event. To determine the active tension (i.e. myosin level) on the new interface formed during the cell neighbor exchange, we assume continuity of myosin concentration at vertices as described in the companion paper [17] and in SI Sec. 4. The active tension is not sufficient to balance the total tension on the new interface, such that passive elements of the cortex (e.g. crosslinkers) are transiently loaded. The resulting passive tension relaxes due to remodeling with timescale $\tau_{\mathrm{p}}$ (see SI Sec. 4 for details). This relaxation causes the elongation of the new interface as it transiently counteracts positive tension feedback and thereby prevents the new interface from immediately re-collapsing after a T1.

### v3

# A Geometric-tension-dynamics Model of Epithelial Convergent Extension

@@ Abstract
Convergent extension of epithelial tissue is a key motif of animal morphogenesis. On a coarse scale, cell motion resembles laminar fluid flow; yet in contrast to a fluid, epithelial cells adhere to each other and maintain the tissue layer under actively generated internal tension. To resolve this apparent paradox, we formulate a model in which tissue flow in the tension-dominated regime occurs through adiabatic remodeling of force balance in the network of adherens junctions. We propose that the slow dynamics within the manifold of force-balanced configurations is driven by positive feedback on myosin-generated cytoskeletal tension. Shifting force balance within a tension network causes active cell rearrangements (T1 transitions) resulting in net tissue deformation oriented by initial tension anisotropy. Strikingly, we find that the total extent of tissue deformation depends on the initial cellular packing order. T1s degrade this order so that tissue flow is self-limiting. We explain these findings by showing that coordination of T1s depends on coherence in local tension configurations, quantified by a geometric order parameter in tension space. Our model reproduces the salient tissue- and cell-scale features of germ band elongation during Drosophila gastrulation, in particular the slowdown of tissue flow after approximately twofold elongation concomitant with a loss of order in tension configurations. This suggests local cell geometry contains morphogenetic information and yields experimentally testable predictions. Defining biologically controlled active tension dynamics on the manifold of force-balanced states may provide a general approach to the description of morphogenetic flow.
NHC contributed equally to this work with FB. NHC carried out the numerical simulations. All authors contributed to conceptualization, analysis, and writing of the manuscript
@@
Shape changes of epithelia during animal development involve major cell rearrangements, often manifested as a “convergent extension” of cell sheets (CE). On the coarse scale, CE resembles the laminar shear flow of an incompressible fluid in the vicinity of a hyperbolic fixed point (see Fig. 1A). Indeed, previous work has combined hydrodynamic equations for the mesoscale cell velocity field with active stress fields to model morphogenetic tissue flow [Oster.etal1983, Streichan.etal2018, Saadaoui.etal2020, Ioratim-Uba.etal2023]. Yet in contrast to a fluid, epithelia are under internally generated tension – as revealed by laser ablation [Kong.etal2019] – and, like solids, maintain their shape against external forces. Tissue flow is achieved through local cell intercalation (T1 neighbor exchange processes, see Fig. 1B) driven by the concerted mechanical activity of individual cells. Cells generate forces via actomyosin contractility in the cortical cytoskeleton at the adherens junctions between cells (Fig. 1C). Moreover, the adherens junctions can remodel through the turnover of their constituent molecules: interfaces in the cell array can change their length and tension independently. This behavior is fundamentally different from (Hookean) springs, where tension and length are related by a constitutive relationship. Instead, one can imagine cellular interfaces as “microscopic muscles” which are actuated by the recruitment and release of myosin motors.
@@
Vertex models generally describe epithelial tissue as a polygonal tiling of cells where the vertex positions are the dynamical variables [Weliky.Oster1990, Farhadifar.etal2007]. The forces that drive the vertex motion are commonly derived from a passive area and perimeter elasticity supplemented with additional active tensions [Collinet.etal2015, Duclut.etal2022, Sknepnek.etal2023]. However, the muscle metaphor for cellular interfaces suggests that active tension is central to the mechanical network underlying an epithelial tissue (Fig. 1D). This network rapidly equilibrates to a force-balanced state [Bonnet.etal2012, Kong.etal2019, Noll.etal2020], stabilized by mechanical feedback loops [Noll.etal2017, Gustafson.etal2022]. In such an active tension network, passive (bulk) elasticity plays a subdominant role. The need for stabilizing feedback loops arises because active tensions are untethered from interface lengths. Indeed, on an abstract level, these feedback loops are not unlike the regulatory mechanisms that control and stabilize skeletal musculature [Byrne1997].
@@
Here, we propose that tissue flow can be understood in terms of adiabatic (quasi-static) remodeling of internal active force balance. Force balance in the cortical tension network defines a manifold of cellular tiling geometries on which tissue deformation unfolds. We propose that dynamics in the force-balance manifold is driven by positive feedback on the cortical tensions. This view is supported by analysis of high-quality live imaging data [Stern.etal2022] from Drosophila gastrulation presented in the companion paper [Brauns.etal2024elife]. Specifically, a geometric active–passive decomposition as well as mutant analysis were used to show that tissue flow is driven by internally generated tension dynamics, rather than external forces. Tension inference has provided evidence for the role of positive tension feedback during active T1 events. Numerical simulations of cell quartets show that such a feedback mechanism is sufficient to drive the T1 process. However, the key question of coordination of T1s across the tissue – required to drive coherent tissue flow – has remained unanswered. To address this question, we develop a model of tissue mechanics in the tension-dominated regime and demonstrate via numerical simulations how positive feedback drives CE. We show that order of the cell packing is necessary for coordinating T1 processes, and hence efficient CE. T1s destroy this order such that the extent of tissue flow is self-limiting. Thereby, our model reproduces the experimentally observed elongation of the germband where the arrest of flow is concomitant with a transition from an ordered to a disordered cell packing [Brauns.etal2024elife].
@@ Methods
Our model is based on two assumptions: (a) on morphogenetic timescales, the forces in the epithelium are approximately balanced, and (b) active cortical tensions (generated by contractile actomyosin along the adherens junctions Fig. 1C) dominate over all other sources of stress. In particular, we assume that adhesion forces between the epithelial layer and its substrate (the fluid yolk and perivitelline fluid [Munster.etal2019] for the Drosophila embryo) are negligible. Hence, all forces must be balanced within the transcellular network of cellular junctions. We model the tissue in the framework of vertex models (see e.g. [Farhadifar.etal2007]) as a polygonal tiling of the plane with tri-cellular vertices $\mathbf{r}_{ijk}$, where each polygon represents a cell $i$ (see Fig 1D). We write the elastic energy differential of this network as
@@
where $\varepsilon$ is a small parameter that separates the dominant scale of active tension and subdominant passive mechanical contributions from bulk and shear elasticity of the cell interior. $E_{\mathcal{C}}(S_{i})$ accounts for the passive elasticity of the cells and will be specified below; $\ell_{ij}=||\mathbf{r}_{ij}||$ and $T_{ij}$ are the length of and tension along the interface between adjacent cells $i$ and $j$. Importantly, in contrast to the standard vertex model where edge tension is defined by a constitutive relation corresponding to a passive Hookean perimeter spring, we take cortical tensions to be controlled independently of the interface lengths. The tension dynamics is described in the next section. The second term in Eq. [1] accounts for the effective in-plane pressure $p$ of the cells that, by maintaining the total surface area (sum over cell areas $A_{i}$), ensures that the tissue as a whole does not collapse. In the Drosophila embryo, the closed epithelial sheet encloses the yolk, which is under pressure, balancing the cortical tensions. Other epithelia are kept under tension via traction forces at the boundary [Kunz.etal2023]. We assume that pressure differences between cells are small and therefore absorb them into $E_{\mathcal{C}}(S_{i})$. In experimental data, this can be verified by inspecting the curvature of cell-cell interfaces: Due to the Laplace law, significant pressure differences would manifest through curved cell-cell interfaces which are not observed in the early Drosophila embryo prior to the onset of cell divisions [Brauns.etal2024elife, Stern.etal2022, Farrell.etal2017, Noll.etal2020].
As noted, we assume a separation of scales between the timescale on which the elastic energy relaxes and the timescale on which the tissue deforms macroscopically. In terms of relaxational dynamics $\gamma\partial_{t}\mathbf{r}_{ijk}=-\frac{\partial E}{\partial\mathbf{r}_{ijk}}$, we consider a relaxation rate – set by the coefficient of friction $\gamma$ – much faster than all other timescales in the system. Quasi-static force balance implies
@@
Solving this equation to zeroth order in $\varepsilon$ yields a force-balance constraint at each vertex: the tension vectors $\mathbf{T}_{ij}=T_{ij}\mathbf{e}_{ij}$ at each vertex must sum to zero and hence form a triangle as illustrated in 1D. Since neighboring vertices share the interface that connects them, the corresponding tension triangles share an edge. Therefore, all tension triangles have to fit together: they form a triangulation that is dual to the cell tiling [Noll.etal2017, Jensen.etal2020, Maxwell1864]. This tension triangulation is a geometric manifestation of global force balance in the tissue, where angles at real-space vertices are complementary to the corresponding angles in the tension triangle (Fig. 1D).
Importantly, fixing the angles at vertices does not fully determine the cell tessellation, i.e. the $\mathbf{r}_{ijk}$: once can change the interface lengths $\ell_{ij}$ while preserving all angles. The resulting isogonal soft modes [^1] A degree-of-freedom count shows there is one isogonal degree of freedom per cell [Noll.etal2017]. Therefore, the isogonal modes can be parametrized by an “isogonal function” that takes a scalar value in each cell. The isogonal displacement of a vertex is defined in terms of the values of this isogonal function in the three adjacent cells (see Eq. [S8] in the SI Appendix). Note that isogonal modes are only soft if they don’t deform the tissue boundary. account for interface length changes under constant tension [Noll.etal2017], which is possible thanks to the turnover of cytoskeletal elements.
Figure 1: A model for tissue mechanics dominated by cortical tension. A Convergent extension (CE) of epithelial tissue by cell intercalations (T1 processes, B). C A single cell from a columnar epithelium with the actomyosin cortex at the adherens junction belt generating tension $T$ along cell-cell interfaces. Gray sphere represents the nucleus. D In force balance, the tensile forces $\mathbf{T}_{ij}$ at each vertex must sum to zero, implying that they form a triangle. The angles in the tension triangulation are complementary to the angles in the cell array, thus linking tension space to physical space. Force balance implies that the tensile force vectors at each vertex form a triangle. E Schematic outline of our quasi-static tissue model with mechanics dominated by actively regulated cortical tensions (see text for details). F Illustration of isogonal deformations of a reference geometry (purple) that is dual to the tension triangulation (red). G Implementation of the model for a symmetric, regular cell array, characterized by one angle $\phi$, determined by the tensions $T_{0},T_{1}$, and two lengths, $\ell_{0}$ and $\ell_{1}$, parametrizing the soft isogonal modes. The contour plot shows the cell shape energy $E_{\mathcal{C}}$ in the incompressible limit where $\ell_{1}$ is determined uniquely by $\ell_{0}$ and $\phi$. Relaxation of the sub-dominant cell shape energy $E_{\mathcal{C}}$ is constrained to the isogonal subspace (white line) determined by tension force balance. For the critical tension ratio $T_{0}/T_{1}=\sqrt{2}$ (corresponding to $\phi_{c}=\pi/2$, black dashed line), the interface length minimizing $E_{\mathcal{C}}$ vanishes (red half-disk). See Movie 1 for an animated version.
The isogonal modes can dilate and shear cells (Fig. 1F). They are soft modes of the leading order elastic energy and thus force us to take into account subleading contributions to arrive at a complete model of tissue mechanics. Cells resist shape distortions due to rigid cell-internal structures such as microtubules, the nucleus [Grosser.etal2021, Kim.etal2022nuclear], and intermediate filaments [Pensalfini.etal2023]. To account for this passive cell elasticity, we propose an energy
@@
where $\mathcal{N}_{i}$ is the set neighbors of cell $i$. This shape energy effectively models the cell interior as a homogeneous elastic material [Wohlrab.etal2024]. (In SI Sec. 5, we compare this elastic energy with the often-used “area–perimeter” elastic energy which we find produces qualitatively different behavior incompatible with experimental observations.) The shape tensor is defined to be invariant under subdivision of interfaces. The reference tensor $S_{0}$ controls the target cell shape and is given by $S_{0}=3\ell_{0}\,\mathbb{I}$ for an isotropic hexagonal cell with side length $\ell_{0}$. We relate cell and tissue elasticity by analyzing the energy spectrum of isogonal modes for a fixed tension triangulation (see SI Sec. 4). The isogonal modes with the lowest energy correspond to large-scale shears and thus provide a linear relationship between the cell and tissue shear moduli (see SI Fig. S13) [^2] As explained in the SI, this analysis only applies for deformations in the bulk of the cell array that leaving the boundary shape unchanged..
@@
For sufficiently small values of the scale-separation parameter $\varepsilon$, minimization of the elastic energy Eq. [1] can be performed in two separate steps: First, force balance of the dominant cortical tensions $T_{ij}$ fixes the angles at vertices thus setting geometric constraints. Second, the sub-leading term $\sum_{i}E_{\mathcal{C}}(S_{i})$ is minimized to fix the remaining isogonal soft modes. Importantly, in this limiting case, the value of $\varepsilon$ is immaterial as long as it is small enough (see SI Sec. 2). Figure 1G and Movie 1 illustrate the minimization of the cell shape energy under the angle constraints imposed by junctional force balance for the minimal setting of a perfectly symmetric cell array. The geometry is characterized by two interface lengths, $\ell_{0}$, $\ell_{1}$, and single angle $\phi$. Incompressibility fixes $\ell_{1}$ as a function of $\ell_{0}$ and $\phi$. We can then plot the cell shape energy $E_{\mathcal{C}}$ in a two-dimensional $\phi,\ell_{0}$ energy landscape. Force balance of the tensions $T_{0}$, $T_{1}$ constrains the angle $\phi=2\arccos(T_{0}/2T_{1}))$ and relaxation of $E_{\mathcal{C}}$ takes place in the isogonal subspace parameterized by $\ell_{0}$ (vertical white line in Fig. 1G), thus fully determining the cell geometry (purple dot). Changing the tensions $T_{0},T_{1}$ shifts the angle constraint and therefore forces the cell shape energy to relax to a new cell geometry. By changing constraints, the dynamics of the tension configuration drives tissue flow. When the tensions reach the critical ratio $T_{0}/T_{1}=\sqrt{2}$ such that $\phi=\pi/2$, the length of the vertical interface, $\ell_{0}$, vanishes, causing a T1 transition as discussed in the companion paper [Brauns.etal2024elife].
@@
Based on previous experiments [Fernandez-Gonzalez.etal2009] and models [Odell.etal1981, Sknepnek.etal2023], we propose a positive feedback mechanism where tension leads to further recruitment of myosin motors and thus further increase in tension. This self-amplifying recruitment is limited by the competition for a limited pool of myosin within each cell. To mimic this effect in a computationally simple way, we constrain tension dynamics to conserve the perimeter of each tension triangle, i.e. the sum of tensions at each vertex $(ijk)$ (see SI Sec. 2 for different local conservation laws). For an individual triangle with tensions $\tilde{T}_{1},\tilde{T}_{2},\tilde{T}_{3}$, we consider the dynamics
@@
where $n$ is an exponent that determines the nonlinearity of the feedback. Note that each cell-cell interface is composed of two actomyosin cortices on its two sides and only one of the two is part of each local pool (see SI Sec. 5 for details). This feedback mechanism has a “winner-takes-all” character, where the longest edge in the tension triangle always outgrows the other two. In our model framework, we can consider a variety of possible local tension dynamics. Below we will also investigate a form of positive tension feedback that saturates, and identify the qualitative features of local tension dynamics key to the tissue dynamics.
@@
Force balance requires that all tension triangles fit together to form a flat triangulation [Noll.etal2017]. The triangulation is parameterized by a set of 2D tension vertex positions $\mathbf{t}_{i}$, so that the tension on edge $(ij)$ is given by $T_{ij}=||\mathbf{t}_{i}-\mathbf{t}_{j}||$. In each iteration of the simulation, the tension vertices $\mathbf{t}_{i}$ are determined by fitting the balanced tensions $T_{ij}$ to the intrinsic tensions $\tilde{T}_{ij}$ using a least squares method. In addition, the intrinsic tensions $\tilde{T}_{ij}$ relax to the balanced tensions $T_{ij}$ with a rate $\tau_{\mathrm{balance}}^{-1}\ll\tau_{\mathrm{T}}^{-1}$ (see SI Sec. 5 for details. All quantifications presented here refer to the flat tensions $T_{ij}$.). This “balancing” of the tension triangulation effectively accounts for small pressure differentials and additional feedback mechanisms (such as the strain rate feedback [Noll.etal2017, Gustafson.etal2022]) which maintain the tension network in a state of force balance. In particular, this ensures that cortical tensions do not lead to a build-up of pressure differentials.
@@
The above dynamics is autonomous in tension space until an edge in the cell tessellation reaches length zero. At this point, a cell neighbor exchange (T1 transition) occurs, corresponding to an edge flip in the tension triangulation. After this topological modification, the tension dynamics continues autonomously again until the next T1 event. To determine the active tension (i.e. myosin level) on the new interface formed during the cell neighbor exchange, we assume continuity of myosin concentration at vertices as described in the companion paper [Brauns.etal2024elife] and in SI Sec. 5. The active tension is not sufficient to balance the total tension on the new interface, such that passive elements of the cortex (e.g. crosslinkers) are transiently loaded. The resulting passive tension relaxes due to remodeling with timescale $\tau_{\mathrm{p}}$ (see SI Sec. 5) [Clement.etal2017]. This relaxation causes the elongation of the new interface, transiently counteracting positive tension feedback and thereby prevents the new interface from immediately re-collapsing after a T1.